OpenStudy (anonymous):

x^2 - 7x +10 = 0 solve, i couldnt get that

OpenStudy (anonymous):

Easiest question ever I'm just stupid.

OpenStudy (anonymous):

yup lol

OpenStudy (anonymous):

\[x^{2}-7x=-10.\] x=2 \[4-14=10\]

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

@omgwut how is x=2?

OpenStudy (anonymous):

2^2=4. 7*2 = 14. 4-14=-10.

OpenStudy (anonymous):

XD XD XD

OpenStudy (anonymous):

@darthjavier bye @omgwut but where did you find the x variable?

OpenStudy (anonymous):

I don't get what you mean.

OpenStudy (anonymous):

how did you know x squared is 2?

OpenStudy (anonymous):

@yomo710 It's a difficult question!! I'm sorry

OpenStudy (anonymous):

x is 2. Not x squared. I use the guess and check method.

OpenStudy (anonymous):

@darthjavier so dont comment?

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

@omgwut im so confused lol

OpenStudy (anonymous):

that works too. x can be 5.

OpenStudy (anonymous):

so how would i write 2 answers? x = (2,5)?

OpenStudy (bibby):

x-5)(x-2)=0 x=2,5

OpenStudy (anonymous):

\(\huge\text{Good God, that's a scary avvy}\)

OpenStudy (anonymous):

\(\Huge\text{ಥ╭╮ಥ}\)